how far has the airplane dropped from the first 600 feet to the airport at -100 feet..
please answer in 10 minutes

Answer by jim_thompson5910(35256)

The plane has dropped 600 - (-100) = 600 + 100 = 700 ft.


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim